Earnings per Share, Price-Earnings Ratio, Dividend Yield The following information was taken from the financial statements of Tolbert Inc. for December 31 of the current fiscal year: Common stock, $20 par value (no change during the year) Preferred $5 stock, $200 par (no change during the year) The net income was $696,000 and the declared dividends on the common stock were $77,500 for the current year. The market price of the common stock is $11.20 per share. For the common stock, determine (a) the earnings per share, (b) the price-earnings ratio, (c) the dividends per share, and (d) the dividend yield. If required, round your answers to two decimal places a. Earnings per Share b. Price-Earnings Ratio c. Dividends per Share d. Dividend Yield $6,200,000 8,000,000
Answer #1
Common shares outstanding 310000 =6200000/20
Dividends on Preferred stock 200000 =(8000000/200)*5
a
Earnings per share =(Net income-Preferred Dividends)/Common Stock outstanding
Earnings per share 1.60 =(696000-200000)/310000
b
Price-Earnings ratio = Market price per share/Earnings per share
Price-Earnings ratio 7 =11.20/1.60
c
Dividends per share 0.25 =77500/310000
d
Dividend yield = Dividend per share/Market price per share
Dividend yield 2.23% =0.25/11.20